OpenStudy (anonymous):

Change the given equation into slope-intercept form. 5x – 2y = -10

OpenStudy (anonymous):

y=5/2x+5

OpenStudy (aroub):

slope intercept form => y=mx+b 5x-2y=-10 -2y=-5x-10 y=5/2x+5

OpenStudy (anonymous):

slope = 5/2 intercept =5

OpenStudy (anonymous):

slope intercept = mx+b the slope of standard form = -a/b, in this case -5/-2 or just 5/2, to find the y intercept of standard form divide the number on the other side of the equals sign by the coefficient of y which is -2 to get 5 so we have y = 5/2x + 5

OpenStudy (anonymous):

-2y=5x - 10